Will you please solve the system by addition or substitution
3x-4y=8
6x-2y=10
thankyou

Answer by funmath(2933)

3x-4y=8
6x-2y=10
-------------------
multiply the 1st equation by -2
-2(3x-4y)=-2(8)
Now your system is:
-6x+8y=-16
6x-2y=10
---------------------------
Add the two equations together:
-6x+8y=-16
+6x-2y=10
______________
0x+6y=-6
6y=-6
6y/6=-6/6
y=-1
---------------------------
Substitute y=-1 into the original first equation and solve for x:
3x-4(-1)=8
3x+4=8
3x+4-4=8-4
3x=4
3x/3=4/3
x=4/3
--------------------------
So the solution of this system is: (4/3,-1)
You can substitute these back into both equations and see if it checks.
3(4/3)-4(-1)=8
4+4=8
8=8 looks like we're right.
6(4/3)-2(-1)=10
8+2=10
10=10 we're absolutely right!